OpenStudy (anonymous):

The index finger of the statue of liberty is 8 ft long. the circumference of the second joint is 3.5 ft. find the approximate volume of the index finger of the statue of liberty. round to the nearest hundreth

OpenStudy (anonymous):

no

OpenStudy (baldymcgee6):

Lets assume that the finger of the statue of liberty is a cylinder.

OpenStudy (anonymous):

so i wouuld use the volume of a cylinder to solve?

OpenStudy (baldymcgee6):

Sure. Do you know the formula for the volume of a cylinder?

OpenStudy (anonymous):

v= pi r2 h

OpenStudy (baldymcgee6):

if you mean V = pi*r^2*h, then yes. You need to use the circumference given to find r.

OpenStudy (anonymous):

so 3.5 is the circumference how would i do that

OpenStudy (baldymcgee6):

well if C = pi*d, then C = pi*2*r, then 3.5 = pi*2*r Solve for r.

OpenStudy (anonymous):

so it would be 21.98

OpenStudy (baldymcgee6):

r = 3.5/(pi*2) = ?

OpenStudy (anonymous):

2.229299363057325

OpenStudy (baldymcgee6):

no

OpenStudy (baldymcgee6):

you must be typing it in wrong.

OpenStudy (baldymcgee6):